# Fast and scalable Wasserstein-1 neural optimal transport solver for single-cell perturbation prediction

### v1

@@ 4.1.2 Wasserstein-2 optimal transport
When the cost function is the square of a distance metric, the optimal transport problem is called Wasserstein-2 optimal transport problem. In this scenario, the Kantorovich duality (3) still holds and it can be transformed into a min-max optimization problem over convex functions [13]:
@@
Based on the elegant and unique solution guaranteed by the Brenier’s theorem and the recently developed input convex neural network (ICNN) [14], OT-ICNN [13] is proposed as a practical Wasserstein-2 optimal transport solver with solid theoretical foundation and CellOT [11] is developed based on it.
@@
When backpropagating through norm-constrained networks, the gradient norms often decrease at each layer as the norm of weight matrix is smaller than 1. This vanishing gradient can be problematic when modeling functions that require preserved input-output gradient norms. For instance, in Wasserstein-1 distance estimation (4), the optimal dual function has a gradient norm of 1 almost everywhere [34]. Previous work has demonstrated that simply constraining the spectral norm of the weight matrices with element-wise, monotonic activation functions (e.g., ReLU) is insufficient to maintain both expressive power and gradient norm preservation during backpropagation [18].
To address this issue, researchers propose enforcing $\mathbf{W}_{l}$ as orthonormal matrices for gradient norm preservation since the norm of an orthonormal matrix is exactly 1 [18]. There are different numerical methods to parameterize orthonormal matrices. Two widely used methods are the Björck orthonormalization algorithm [35, 18] and the Cayley transform [36, 37]. The Björck algorithm iteratively computes the closest orthonormal matrix by applying a Taylor expansion of the polar decomposition, and the Cayley transform can map a skew-symmetric matrix to an orthogonal matrix. In practice, the high-order Björck algorithm with enough iterations leads to a high-precision approximation but requires more computation, while the Cayley transform performs faster but with less precision.
@@
1-Lipschitz activation functions are common in neural networks as long as the slope is less than or equal to 1 (e.g., ReLU). However, to maintain the expressive power of a 1-Lipschitz neural network, the activation function should also preserve the gradient norm [18]. The GroupSort activation function was proposed to address this need by preserving the gradient norm. It operates by first splitting the features into groups and then sorting the features within each group to produce the output. The gradient norm-preserving property arises from the fact that its Jacobian is a permutation matrix, and permutation matrices preserve every vector norm [18]. Moreover, previous theoretical analyses have shown that a GroupSort neural network can represent any Lipschitz continuous piecewise linear function, making them well-suited for approximating Lipschitz continuous functions [19]. Therefore, we employ the GroupSort function with an orthonormalized linear layer to build the 1-Lipschitz neural network.
@@ 4.2.1 $W_{1}$ optimal transport solver
At the first step, we aim to parameterize the 1-Lipschitz function in (4) as 1-Lipschitz neural network to maximize the dual problem. The training objective can be described as:
@@
Next, after having a fitted Kantorovich potential, we need to learn the transportation step size based on the direction of $\nabla f_{\theta}(x)$. However, there is not any principled method to solve this problem since the Kantorovich dual only weaken the primal problem rather than solving it in the Wasserstein-1 settings [23, 22]. Most recently, a method is proposed to learn a fixed step size for all samples from the data [38], but fixed step size is not accurate and will distort the desired alignment between transported and target distribution. Here, we propose to parameterize the step size function as a neural network and train it under the generative adversarial networks (GAN) paradigm to learn a sample-specific transportation. Specifically, we define a step size function $\eta_{\omega}:\mathbb{R}^{d}\mapsto\mathbb{R}^{+}$ parameterized by $\phi$, which takes the input sample $x$ and outputs a positive step size. If we fix the parameters of the learned potential function, then the transport map $T_{\omega}$ is then defined as:
@@
To train the step size function, we employ a discriminator network $D_{\xi}:\mathbb{R}^{d}\mapsto[0,1]$ parameterized by $\xi$, which aims to distinguish between the transported samples and the target distribution samples. The training objectives for the step size function and the discriminator are as follows:
@@
The $W_{2}$ optimal transport solver typically relies on the min-max optimization of the Kantorovich dual and converges slowly [15]. Although our method avoids this step when learning the Kantorovich potential, we still employ a GAN to learn the step size function, which inherently involves a min-max optimization problem. However, we emphasize that solving the $W_{1}$ dual is beneficial as it simplifies the primal problem and provides direction for transportation. With this directional guidance, the GAN optimization is significantly simplified, as the generator only needs to produce a scalar step size. Thus, we believe the key advantage of our method lies in breaking down the complex optimal transport primal problem into two simpler optimization tasks.
@@ 4.2.2 Baselines
To benchmark our method’s performance, we include four baseline methods for comparison. The first baseline is the identity transformation, where no transformation is applied to the source distribution data. The second baseline is the observed target distribution, which uses the observed target distribution data from the training set for evaluation. The third baseline is scGen [7], which computes the displacement between the means of the source and target distributions and applies this displacement for transportation. Following the benchmark procedure in CellOT [11], we implement an autoencoder for scGen. Notably, we balance the training data by equalizing the number of source and target samples for scGen to avoid bias; otherwise, we observe that the performance of scGen would be worse than the identity transformation. Lastly, we implement the $W_{2}$ optimal transport solver, following the implementation in OT-ICNN [13].
@@ 4.2.3 Hyperparameters and training details
We employ the hyperparameters used in CellOT [11] for both the autoencoder and the $W_{2}$ Kantorovich potential training. Specifically, the hidden layers for the encoder and decoder are set to [512, 512] for the single-cell RNA-seq dataset (sciplex3) [6] and [32, 32] for the single-cell imaging dataset (4i) [4]. The latent dimension is set to 50 for the sciplex3 and 8 for the 4i. The autoencoder is trained for 250,000 iterations using the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-5}$. The checkpoint with the best validation training loss will be saved and used as the final model. The $W_{2}$ Kantorovich potential is parameterized as an input convex neural network with hidden layers of size [64, 64, 64, 64]. The $W_{2}$ objective (7) is optimized over 100,000 iterations, with 10 inner iterations per optimization step, using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and betas (0.5, 0.9) [13].
For a fair comparison, the $W_{1}$ potential function employs the same hidden layer architecture as the $W_{2}$ model. This potential function is parameterized as a 1-Lipschitz network using the Cayley transform for fast orthonormalization approximation. According to previous work [19], which shows that a larger grouping size could lead to a more expressive network, we test the effect of grouping size on synthetic data and observe that a larger grouping size does not always lead to better performance within a fixed number of training iterations. The explanation for this phenomenon is that if the grouping size equals the hidden units, the activation would degenerate into an identity transformation. Based on this observation, we set the grouping size moderately larger than 1, choosing 4 to achieve a more expressive network and to avoid the degeneration issue. This network is trained for 10,000 iterations using the Adam optimizer with betas (0.5, 0.5), along with a cosine annealing scheduler. The learning rate starts at $1\times 10^{-2}$ and is gradually adjusted to $1\times 10^{-4}$. To learn the step size function, both the discriminator and the step size function use hidden layers of size [64, 64, 64, 64] and are optimized for 10,000 iterations with the Adam optimizer at learning rate of $1\times 10^{-4}$.
@@ 4.3.1 Synthetic toy datasets
To verify that the transportation map is “monotonic”, we generate the “bookshelf” and “circles” datasets. For the “bookshelf” dataset, we created two narrow vertical strips. The source data consists of points where the x-values are uniformly distributed between 0 and 1, and the y-values are normally distributed around zero with a standard deviation of 0.001. The target data is similarly generated, but with x-values uniformly distributed between 2 and 3. This design results in two parallel strips resembling books on a shelf. For the “circles” dataset, we used the scikit-learn make_circles function to generate concentric circles with added noise. Besides evaluating the “monotonic” property, we evaluate the nonlinearity modeling ability with the “swiss roll” dataset, which is generated using the make_swiss_roll function. The source data is drawn from a standard 2D Gaussian distribution. The target data consists of two dimensions extracted from the 3D swiss roll and scaled appropriately. For a more complex nonlinear example, we generate the “moons” dataset utilizing the make_moons function to create two interleaving half-circles with added noise. The source and target data are composed of points from each of the two moons, respectively.
@@ 4.3.2 Single-cell datasets
We employ the “sciplex3” [6] and “4i” [4] datasets to evaluate our method on real single-cell perturbation datasets. The “4i” dataset consists of 37 perturbations along with a control group. It contains 48 cell imaging features after de-duplication and is less sparse compared to conventional scRNA-seq datasets. Feature selection follows the processing procedure of CellOT, which filters out the sum of the protein intensity features and retains the mean. For the “sciplex3” scRNA-seq dataset, we normalize and log-transform the data, retaining the top 1,000 highly variable features. The processed sciplex3 data is provided by the original CellOT publication.
@@ 4.4 Evaluation metrics
We evaluate the effectiveness of different methods using three key metrics: the average correlation coefficient $r^{2}$ of feature means, the $l_{2}$ feature mean distance, and the maximum mean discrepancy (MMD). The $l_{2}$ feature means measure the Euclidean distance between the means of observed and predicted distributions, while the $r^{2}$ correlation coefficient quantifies the similarity between the observed and predicted means. However, relying solely on metrics that evaluate feature means can be limiting in cases where the target distribution exhibits multimodality or other complex characteristics. To effectively evaluate the performance, we employ MMD, a distributional distance metric sensitive to higher-order moments, which helps us capture discrepancies between predicted and real distributions beyond simple means [26]. Following the implementation in CellOT, MMD is computed using an RBF kernel and averaged over several length scales to ensure robustness across varying scales of the data [26].

### v2

@@ 2.2.2 Wasserstein-2 optimal transport
When the cost function is the square of a distance metric, the optimal transport problem is called Wasserstein-2 optimal transport problem. In this scenario, the Kantorovich duality (3) still holds and it can be transformed into a min-max optimization problem over convex functions (Makkuva et al., 2020):
@@
$-\int_{\mathcal{X}}\langle x,\nabla\psi(x)\rangle-\phi(\nabla\psi(x))\ d\mu(x)+C_{\mathcal{\mu},\mathcal{\nu}}$ (7)
@@
Based on the elegant and unique solution guaranteed by the Brenier’s theorem and the recently developed input convex neural network (ICNN) (Amos et al., 2017), OT-ICNN (Makkuva et al., 2020) is proposed as a practical Wasserstein-2 optimal transport solver with solid theoretical foundation and CellOT (Bunne et al., 2023) is developed based on it.
@@
When backpropagating through norm-constrained networks, the gradient norms often decrease at each layer as the norm of weight matrix is smaller than 1. This vanishing gradient can be problematic when modeling functions that require preserved input-output gradient norms. For instance, in Wasserstein-1 distance estimation (4), the optimal dual function has a gradient norm of 1 almost everywhere (Gulrajani et al., 2017). Previous work has demonstrated that simply constraining the spectral norm of the weight matrices with element-wise, monotonic activation functions (e.g., ReLU) is insufficient to maintain both expressive power and gradient norm preservation during backpropagation (Anil et al., 2019).
To address this issue, researchers propose enforcing $\mathbf{W}_{l}$ as orthonormal matrices for gradient norm preservation since the norm of an orthonormal matrix is exactly 1 (Anil et al., 2019). There are different numerical methods to parameterize orthonormal matrices. Two widely used methods are the Björck orthonormalization algorithm (Björck and Bowie, 1971; Anil et al., 2019) and the Cayley transform (Cayley, 1846; Trockman and Kolter, 2021). The Björck algorithm iteratively computes the closest orthonormal matrix by applying a Taylor expansion of the polar decomposition, and the Cayley transform can map a skew-symmetric matrix to an orthogonal matrix. In practice, the high-order Björck algorithm with enough iterations leads to a high-precision approximation but requires more computation, while the Cayley transform performs faster but with less precision.
@@
1-Lipschitz activation functions are common in neural networks as long as the slope is less than or equal to 1 (e.g., ReLU). However, to maintain the expressive power of a 1-Lipschitz neural network, the activation function should also preserve the gradient norm (Anil et al., 2019). The GroupSort activation function was proposed to address this need by preserving the gradient norm. It operates by first splitting the features into groups and then sorting the features within each group to produce the output. The gradient norm-preserving property arises from the fact that its Jacobian is a permutation matrix, and permutation matrices preserve every vector norm (Anil et al., 2019). Moreover, previous theoretical analyses have shown that a GroupSort neural network can represent any Lipschitz continuous piecewise linear function, making them well-suited for approximating Lipschitz continuous functions (Tanielian and Biau, 2021). Therefore, we employ the GroupSort function with an orthonormalized linear layer to build the 1-Lipschitz neural network.
@@ 2.3.1 $W_{1}$ optimal transport solver
At the first step, we aim to parameterize the 1-Lipschitz function in (4) as 1-Lipschitz neural network to maximize the dual problem (Fig. 1b). The training objective can be described as:
@@
Next, after having a fitted Kantorovich potential, we need to learn the transportation step size based on the direction of $\nabla f_{\theta}(x)$. However, there is not any principled method to solve this problem since the Kantorovich dual only weaken the primal problem rather than solving it in the Wasserstein-1 settings (Caffarelli et al., 2002; Trudinger and Wang, 2001). Most recently, a method is proposed to learn a fixed step size for all samples from the data (Milne et al., 2022), but fixed step size is not accurate and will distort the desired alignment between transported and target distribution. Here, we propose to parameterize the step size function as a neural network and train it under the generative adversarial networks (GAN) paradigm to learn a sample-specific transportation. Specifically, we define a step size function $\eta_{\omega}:\mathbb{R}^{d}\mapsto\mathbb{R}^{+}$ parameterized by $\phi$, which takes the input sample $x$ and outputs a positive step size. If we fix the parameters of the learned potential function, then the transport map $T_{\omega}$ is then defined as (Fig. 1b):
@@
To train the step size function, we employ a discriminator network $D_{\xi}:\mathbb{R}^{d}\mapsto[0,1]$ parameterized by $\xi$, which aims to distinguish between the transported samples and the target distribution samples (Fig. 1b). The training objectives for the step size function and the discriminator are as follows:
@@
The $W_{2}$ optimal transport solver typically relies on the min-max optimization of the Kantorovich dual and converges slowly (Korotin et al., 2021). Although our method avoids this step when learning the Kantorovich potential, we still employ a GAN to learn the step size function, which inherently involves a min-max optimization problem. However, we emphasize that solving the $W_{1}$ dual is beneficial as it simplifies the primal problem and provides direction for transportation. With this directional guidance, the GAN optimization is significantly simplified, as the generator only needs to produce a scalar step size. Thus, we believe the key advantage of our method lies in breaking down the complex optimal transport primal problem into two simpler optimization tasks.
@@ 2.3.2 Baselines
To benchmark our method’s performance, we include four baseline methods for comparison. The first baseline is the identity transformation, where no transformation is applied to the source distribution data. The second baseline is the observed target distribution, which uses the observed target distribution data from the training set for evaluation. The third baseline is scGen (Lotfollahi et al., 2019), which computes the displacement between the means of the source and target distributions and applies this displacement for transportation. Following the benchmark procedure in CellOT (Bunne et al., 2023), we implement an autoencoder for scGen. Notably, we balance the training data by equalizing the number of source and target samples for scGen to avoid bias; otherwise, we observe that the performance of scGen would be worse than the identity transformation. Next, we implement the $W_{2}$ optimal transport solver, following the implementation in OT-ICNN (Makkuva et al., 2020) and CellOT (Bunne et al., 2023). Of note, CellOT is just a new PyTorch implementation of OT-ICNN, showing OT’s application on single-cell datasets. Hence, we use the name “$W_{2}$ OT” in our manuscript to denote these two methods. Additionally, we include CPA (Lotfollahi et al., 2023) and biolord (Piran et al., 2024) for the performance reference.
@@ 2.3.3 Hyperparameters and training details
We employ the hyperparameters used in CellOT (Bunne et al., 2023) for both the autoencoder and the $W_{2}$ Kantorovich potential training. Specifically, the hidden layers for the encoder and decoder are set to [512, 512] for the single-cell RNA-seq dataset (sciplex3) (Srivatsan et al., 2020) and [32, 32] for the single-cell imaging dataset (4i) (Gut et al., 2018). The latent dimension is set to 50 for the sciplex3 and 8 for the 4i. The autoencoder is trained for 250,000 iterations using the Adam optimizer with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-5}$. The checkpoint with the best validation training loss will be saved and used as the final model. The $W_{2}$ Kantorovich potential is parameterized as an input convex neural network with hidden layers of size [64, 64, 64, 64]. The $W_{2}$ objective (2.2.2) is optimized over 100,000 iterations, with 10 inner iterations per optimization step, using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and betas (0.5, 0.9) (Makkuva et al., 2020).
For a fair comparison, the $W_{1}$ potential function employs the same hidden layer architecture as the $W_{2}$ model. This potential function is parameterized as a 1-Lipschitz network using the Cayley transform for fast orthonormalization approximation. According to previous work (Tanielian and Biau, 2021), which shows that a larger grouping size could lead to a more expressive network, we test the effect of grouping size on synthetic data (Korotin et al., 2022) and observe that a larger grouping size does not always lead to better performance within a fixed number of training iterations (Supplementary Fig. S1). The explanation for this phenomenon is that if the grouping size equals to the hidden units, the activation would degenerate into an identity transformation. Based on this observation, we set the grouping size moderately larger than 1, choosing 4 to achieve a more expressive network and to avoid the degeneration issue. This network is trained for 10,000 iterations using the Adam optimizer with betas (0.5, 0.5), along with a cosine annealing scheduler. The learning rate starts at $1\times 10^{-2}$ and is gradually adjusted to $1\times 10^{-4}$. To learn the step size function, both the discriminator and the step size function use hidden layers of size [64, 64, 64, 64] and are optimized for 10,000 iterations with the Adam optimizer at learning rate of $1\times 10^{-4}$.
For the CPA model, we use the default hyperparameters provided in its repository. In our experiments, we find that these default settings perform reasonably well on the “sciplex3” dataset but much worse on the “4i” dataset. As there are too many hyperparameters, we then only sweep the penalty parameters and adversary steps on ”4i” dataset and choose to set penalty_adversary=100, reg_adversary=0, and adversary_steps=6 for “4i” dataset. For the biolord model, we only tune the loss penalty hyperparameters using a random perturbation condition of the “sciplex3” dataset (1 out of 9), while keeping all other hyperparameters fixed to their reported best values. Based on the search results, we set unknown_attribute_penalty=100 and reconstruction_penalty=100 for all experiments presented in the paper.
@@ 2.4.1 Synthetic toy datasets
We first use the synthetic data from the Wasserstein-1 solver benchmark (Korotin et al., 2022) to choose the appropriate group size by testing the group sort network on synthetic data with different dimensions and different number of funnels (Supplementary Fig. S1). Next, to verify that the transportation map is “monotonic”, we generate the “bookshelf” and “circles” datasets. For the “bookshelf” dataset, we created two narrow vertical strips. The source data consists of points where the x-values are uniformly distributed between 0 and 1, and the y-values are normally distributed around zero with a standard deviation of 0.001. The target data is similarly generated, but with x-values uniformly distributed between 2 and 3. This design results in two parallel strips resembling books on a shelf. For the “circles” dataset, we used the scikit-learn make_circles function to generate concentric circles with added noise. Besides evaluating the “monotonic” property, we evaluate the nonlinearity modeling ability with the “swiss roll” dataset, which is generated using the make_swiss_roll function. The source data is drawn from a standard 2D Gaussian distribution. The target data consists of two dimensions extracted from the 3D swiss roll and scaled appropriately. For a more complex nonlinear example, we generate the “moons” dataset utilizing the make_moons function to create two interleaving half-circles with added noise. The source and target data are composed of points from each of the two moons, respectively.
@@ 2.4.2 Single-cell datasets
We employ the “sciplex3” (Srivatsan et al., 2020) and “4i” (Gut et al., 2018) datasets to evaluate our method on real single-cell perturbation datasets. The “4i” dataset consists of 37 perturbations along with a control group. It contains 48 cell imaging features after de-duplication and is less sparse compared to conventional scRNA-seq datasets. Feature selection follows the processing procedure of CellOT, which filters out the sum of the protein intensity features and retains the mean. For the “sciplex3” scRNA-seq dataset, we normalize and log-transform the data, retaining the top 1,000 highly variable features. The processed sciplex3 data is provided by the original CellOT publication.
@@ 2.5 Evaluation metrics
We evaluate the effectiveness of different methods using three key metrics: the average correlation coefficient $r^{2}$ of feature means, the $l_{2}$ feature mean distance, and the maximum mean discrepancy (MMD). The $l_{2}$ feature means measure the Euclidean distance between the means of observed and predicted distributions, while the $r^{2}$ correlation coefficient quantifies the similarity between the observed and predicted means. However, relying solely on metrics that evaluate feature means can be limiting in cases where the target distribution exhibits multimodality or other complex characteristics. To effectively evaluate the performance, we employ MMD, a distributional distance metric sensitive to higher-order moments, which helps us capture discrepancies between predicted and real distributions beyond simple means (Gretton et al., 2012). Following the implementation in CellOT, MMD is computed using an RBF kernel and averaged over several length scales to ensure robustness across varying scales of the data (Gretton et al., 2012).